OpenStudy (anonymous):

solve the proportion g/56 = 2/8 help!

OpenStudy (anonymous):

what is 56/8?

OpenStudy (anonymous):

fraction

OpenStudy (anonymous):

How many times does 8 go into 56?

OpenStudy (anonymous):

7

OpenStudy (michele_laino):

we have to apply the fundamental property of proportion, so we can write: \[\Large g = \frac{{56 \times 2}}{8} = ...?\]

OpenStudy (anonymous):

7*2=?

OpenStudy (anonymous):

I really need help with this u guys idk now much

OpenStudy (anonymous):

7 x 2 = ? g = answer

OpenStudy (anonymous):

7 x 2 = 14

OpenStudy (anonymous):

There you go

OpenStudy (anonymous):

that's the answer?

OpenStudy (anonymous):

14/56 = 2/8

OpenStudy (anonymous):

That is the ansewr

OpenStudy (anonymous):

oh ya thank you!

OpenStudy (anonymous):

Do you understand how you worked that out?

OpenStudy (anonymous):

not really I suck at math I just go o google A LOT OF THE TIMES!

OpenStudy (kendricklamar2014):

\[\frac{ g }{ 56 } = \frac{ 2 }{ 8 }\] Cross Multiply: \[g*(8)=(2)*(56)\] \[= 8g=112\] Divide both sides by 8: \[\frac{ 8g }{ 8 } = \frac{ 112 }{ 8 }\]

OpenStudy (kendricklamar2014):

\[\frac{ 8g }{ 8 } = g\] \[\frac{ 112 }{ 8 } = ?\] @ilyvowls2004